OpenStudy (anonymous):

Given the equation y − 3 =1/2 (x + 6) in point-slope form, identify the equation of the same line in slope-intercept form

OpenStudy (mathmate):

Slope intercept form is y=mx+b To proceed, expand both sides of the original equation, and isolate y on the left hand side. Collect terms and simplify on the right.

OpenStudy (anonymous):

y − 3 =1/2 (x + 6), solve for y y= 1/2 x +3+3 y=1/2 x +6